yardstick ninja 2. Life has been good to you so far. Now you have a job, and, youre thinking, life is even better now that yLife has been good to you so far. Now you have a job, and, you're thinking, life is even better now that you have a paycheck! However, your job has been trying, and a little bit boring. So, you'd like to take up some time developing an algorithm to make the job even easier than it already is. Here's you job: You are the Measure Minister. People in your town, being too stupid to use a yardstick , bring to you all their stuff to measure. Everything is measured in exact inches. (The foregoing fact makes this problem simple, albeit not realistic. Thus, for this problem, you can assume no one will ever bring you anything measuring 3.5 inches, or 78.23 inches, or 9.01 inches, or ...) Your only tool-o-measurement is a piece of string exactly 8 inches long. Write an algorithm that you (or your replacement) may use to measure in the most efficient way all the stuff that the townspeople might bring to you to measure. Remember, they can bring you little things or big things...or even really big things! Remember, your measures will always be to the nearest inch, and the string you use is exactly 8 inches long. yardştick ninja.
Answer #1

All the things can be measured to the nearest inch if we can fold the string.

One way is to fold the string in half 3 times to make it exactly 1 inch long(8/2 = 4, 4/2 = 2, 2/2 = 1). This way we can use the one inch string to measure the objects.

Other way is to measure the objects using 8 inch string, if the remaining part of the object is less than 8 inches long, fold the string and then measure. For eg, to measure 25 inches, first use 8 inches string 3 times(we get 24) and then fold the string 3 times to get 1 inch string and use it to measure the remaining part.

Both the methods will get the same answer. The first one will take longer, the second one will be faster.

Hope this helps.

Please rate the answer if you like it.

Do leave a comment.An suggestion/query is much appreciated.